# Quasinormal spectra of a wormhole family: overtone features and a parameter-controlled redshift

### v1

@@ Abstract
Finally, we show how an effect somewhat similar to the so-called ‘environment induced redshift’ of QNMs, introduced recently (arXiv:2412.18651), may be obtained for spacetimes with non-constant $g_{00}$, via an appropriate tuning of available metric parameters which systematically modify the shapes of the effective potentials arising in the perturbation equations.
@@ I Introduction and summary
The entire scenario was largely theoretical, with very little effort being given to even suggest possible observational consequences. A large part of the efforts focused on how to avoid energy condition violations Visser et al. (2003); Maeda and Nozawa (2008), and modified/alternate theories of gravity Bhawal and Kar (1992); Kanti et al. (2012); Hochberg (1990); Agnese and La Camera (1995); Shaikh and Kar (2016); Kumar et al. (2024); Radhakrishnan et al. (2024); Dehghani and Dayyani (2009); Sadatian and Hosseini (2024) seemed to be a sort of way out. It was noticed that in a modified theory of gravity, the purely geometric violation of the timelike or null convergence condition does not necessarily imply a violation of the weak, null or dominant energy conditions. Hence, energy conditions, postulated as physical requirements on matter, may hold even though the convergence conditions may be violated, as required for a wormhole to exist. Thus, wormholes seemed more viable in various modified theories of gravity.
@@
Given the above fact that wormholes could be around if any modified theory of gravity is acceptable, even in a restricted domain of the theory parameters, it is obvious that one must find observational signatures. Gravitational lensing of wormholes Cramer et al. (1995); Dey and Sen (2008); Shaikh et al. (2019); Godani and Samanta (2021); Nandi et al. (2006); Eiroa et al. (2001); Izmailov et al. (2019) was one of the first among the observationally relevant works pursued by researchers. It still continues as an important possible signature.
@@
Over the last decade or so, astronomy and astrophysics has reached a different level of observational relevance through the emergence of a large amount of data and results in gravitational wave astronomy (LIGO-VIRGO-KAGRA collaboration Abbott et al. (2023)) and infrared imaging of astrophysical objects (the Event Horizon Telescope (EHT) Akiyama et al. (2019, 2022)). The results obtained tally well with General Relativity predictions, leaving little room for modified gravity. However, there does remain a small but definite possibility of accommodating a modified gravity solution or a sort of exoticity within GR (eg. the Bronnikov-Ellis geometry). It is this possibility which has driven some researchers (including us) to construct models, do calculations with the aim of looking for signatures of wormholes in available data. With more observed events and data in the future, one can always hope for further confirmation of GR or, perhaps, some acceptance of a modification of GR, of course, within a specific domain of parameter values.
@@
One of the central observationally important arenas in GW astrophysics, which has been flooded, in the last decade, by a huge amount of theoretical work by various groups, is the study of quasinormal modes Vishveshwara (1970) in perturbation theory (see Berti et al. (2025) for a very recent review). Scalar and gravitational perturbations have been worked out for diverse black hole Berti ; Berti et al. (2009); Cardoso and Lemos (2001); Nollert (1999); Konoplya (2002, 2003); Rincón and Santos (2020); Blázquez-Salcedo et al. (2018a) and wormhole Bueno et al. (2018); Konoplya and Molina (2005); Kim (2008); Alfaro et al. (2024); Jirón et al. (2025) spacetimes in GR as well as in modified gravity theories. Any such study on a spacetime not among the conventional solutions of GR (eg. Schwarzschild or Kerr) is referred to as a study on a ‘mimicker’. More specifically, wormholes constitute a prominent family among such ‘black hole mimickers’ Abdikamalov et al. (2019); Konoplya et al. (2019); Urbano and Veermäe (2019); Narzilloev et al. (2020); Olivares et al. (2020); Shaikh et al. (2022); Shao et al. (2021); Siemonsen (2024); Pani et al. (2010); Cardoso and Pani (2019); Konoplya and Zhidenko (2016); Glampedakis and Pappas (2018); De Simone et al. (2025). Collectively, such mimickers are also referred to as ‘exotic compact objects’. The main idea here is that these spacetimes could be among possible stable end-states in a post-merger scenario Sanchis-Gual et al. (2019); Cardoso et al. (2016); Krishnendu et al. (2017). Whether this is a feasible option at all, will be evident as more observational results emerge.
@@
$\bullet$ Most of the work on QNMs of the family of spacetimes studied here is focused on the fundamental QNMs, which surely have the largest contribution (amplitude-wise) in a time domain profile. Despite this, with possible detection Capano et al. (2023) (though disputed Cotesta et al. (2022)) of overtones and the possible detectability of overtones in LISA Oshita and Tsuna (2023), it is always important to track down the contributions of the so-called ‘overtones’. This, indeed, has been a line of thought of late, with varied motivations. Researchers were keen on finding overtones from the onset of QNM analysis. Chandrasekhar and Detweiler Chandrasekhar and Detweiler (1975) computed the fundamental mode along with a couple of overtones for the Schwarzschild metric. Later, the QNM spectra with overtones for RN Ferrari and Mashhoon (1984), Kerr Detweiler (1980); Leaver (1985), Kerr-Newman Berti and Kokkotas (2005), regular blackhole Konoplya et al. (2022); Konoplya and Zhidenko (2024); Zhang et al. (2024), parametrized black hole Hirano et al. (2024) metrics were computed. The authors of Kim (2008); Batic and Dutykh (2024); Alfaro et al. (2024); Ávalos and Contreras (2022); Jirón et al. (2025); González et al. (2022) computed QNM overtones for various wormhole geometries.
@@
$\bullet$ It has been found (theoretically) that black hole quasinormal modes, when obtained for a scenario where the black hole is in an environment (e.g. a halo), exhibit a ‘redshift’ w.r.t. the increase of the environment parameter Barausse et al. (2014); Pezzella et al. (2025); Liu et al. (2025).
We show here, for the QNMs and overtones found by us for the wormhole family, that there is indeed a ‘redshift’ when we modify the so-called effective potential by tuning a metric parameter appropriately. In this context, we also provide additional examples of a similar redshift in the simplest of examples–eg. the square barrier or the Poschl-Teller potential.
Our article, which is largely a sequel to Dutta Roy et al. (2020), is organised as follows. In the next section, we provide a summary of the wormhole family we study in this article. Section III quickly outlines the methods adopted in finding the QNMs–more specifically, which method is better suited for finding overtones is emphasised and discussed in more detail here. In Section IV, we summarise the numerical results obtained with explicit plots. Section V discusses the inferences which may be drawn from our studies on the QNMs and overtones. The above-mentioned ‘redshift’ effect is discussed extensively in Section VI. Finally, in Section VII, we conclude with remarks on possible future work. The appendix contains the detailed numerical values of the QNMs, thereby supporting the plots shown in Section IV.
@@ II A brief review on the family of wormhole geometries
One may also rewrite the line element using the proper radial distance, i.e $r^{2}=b_{0}^{2}+l^{2}$ (where $-\infty\leq l\leq\infty$). In this representation, we have
@@
We may visualise the above line element as representing two asymptotically flat regions ($l\rightarrow\pm\infty$) connected by a throat/bridge of radius $b_{0}$. Note that the radial coordinate $r$ runs from $b_{0}$ to $\infty$ whereas $l$ runs from $\infty$- ∞ to $\infty$ with $l=0$ corresponding to $r=b_{0}$.
@@ II.1 The ultra-static wormhole family
Here, $l$ ([,])absent($\in$[$\infty$,$\infty$])( ∈ [ - ∞ , ∞ ] ) is the tortoise coordinate
@@ II.2 Bronnikov–Ellis wormhole with asymptotic mass
Till now, we have discussed the ultrastatic, symmetric and massless Bronnikov–Ellis geometry and its generalisation. However, the authors of Bronnikov (1973), Ellis (1973) (see also Blázquez-Salcedo et al. (2018b)) have proposed a metric for a massive, static, Bronnikov–Ellis wormhole for which the line element is of the following form:
@@
$f(l)=\frac{C}{b_{0}}\left(arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$
@@
For $M=0$, we recover the symmetric generalised Bronnikov–Ellis wormhole, while any non-zero value of $M$ will introduce asymmetry in the wormhole shape. The QNMs for the $n=2$, $M\neq 0$ geometry will be discussed in Section VI.3.
@@
However, in the case of $M\neq 0$ and $n>2$, one can find that all three conditions are satisfied in certain regions if the $|M|$ exceeds a specific value. For different geometries (different values of $n$), the critical value of $M$ and the acceptable regions (corresponding to a fixed value of $M$) are different.
@@ II.3 Bronnikov–Ellis geometry with a chosen redshift function:
It is also possible to choose a suitable redshift function ($-g_{00}$ never zero) in an ad-hoc way while retaining the Bronnikov–Ellis $n=2$ character unaltered in the spacelike slice. The ensuing consequences for required matter and energy conditions can thereafter be looked at. As an example, let us consider the line element studied in Alfaro et al. (2024) and given by,
@@
Note that the energy density is not affected by $M$, though the pressures change. The Weak Energy Condition is violated for all $r>b_{0}$ since the energy density is negative. The Null Energy Condition is also violated with complementary behaviour for the $\rho+\tau$ and $\rho+p$ inequalities w.r.t. increase in $M$ value. As $M$ increases the region of violation (near $r=b_{0}$) of the $\rho+\tau$ inequality shrinks whereas the region of violation (away from $r=b_{0}$) of the $\rho+p$ inequality becomes larger. We have plotted $\rho+\tau$ and $\rho+p$ as a function of the radial coordinate $r$, for two values of $M$ (Fig. 2), in order to show this feature in the NEC violations.
Putting $M=0$, we get back the original Bronnikov–Ellis wormhole. We will discuss the effective potential and the $\omega_{QNM}$ for this spacetime in a later section (VI.4).
@@ II.4 Effective potentials for scalar and axial perturbation
The original Bronnikov–Ellis wormhole ($n=2$) is a solution to the Einstein equations coupled to a phantom scalar field. The $n>2$ geometries also have contributions from the phantom field. The perturbations in this phantom scalar field ($\Phi$) can give rise to scalar quasinormal modes. For this, we may write the scalar field as $\phi+\Phi$.
@@
$\frac{d^{2}\psi}{dl_{*}^{2}}+(\omega^{2}-V_{sc}(l))\psi=0\hskip 56.9055ptwhere\hskip 28.45274pt\frac{dl_{*}}{dl}=e^{-f}$ (16)
@@
III Methods used to find the QNMs
III.1 Time domain and Prony fit
The time domain profile represents the temporal evolution of the field $\psi$ in a fixed background geometry. It was originally formulated in Gundlach et al. (1994). In the previous section, we derived the master wave equation, which may also be rewritten without the separability ansatz for the field. Thus, the equation, using the new coordinates $u$ and $v$, such that $du=dt-dx,\hskip 5.69046ptdv=dt+dx$, reads Konoplya and Zhidenko (2011):
$\frac{\partial^{2}\psi}{\partial u\partial v}+V(u,v)\psi=0.$ (20)
where $S=(u,v)$, $W=(u+h,v)$, $E=(u,v+h)$, and $N=(u+h,v+h)$ and $h$ is the integration step size for both $u$ and $v$. Eqn. (21) enables us to compute $\psi$ iteratively on the $u-v$ grid based on the initial data specified on null surfaces $u=u_{0}$ and $v=v_{0}$. The values of $\psi(N)$, thus calculated, give us the time-domain profile for the given potential.
The time-domain profile of $\psi$ is then fitted using the Prony method $\psi(t)\approx\sum_{j=1}^{p}C_{j}e^{i\omega_{j}t}$. Here $\omega_{j}$ are the quasinormal frequencies and $C_{j}$ are the corresponding amplitudes. We assume that the ringdown starts at $t=0$ and ends at $t=N_{1}h$, and approximate the profile data as a sum of $p\hskip 2.84544pt(\leq\frac{N_{1}+1}{2})$ terms in the following way:
$x_{n}\approx\psi(nh)=\sum_{j=1}^{p}C_{j}z_{j}^{n},\hskip 56.9055ptz_{j}=e^{-i\omega_{j}nh}$ (22)
where $a_{0}=1$. This leads to the matrix equation:
$\begin{bmatrix}x_{p-1}&x_{p-2}&\ldots&x_{0}\\ x_{p}&x_{p-1}&\ldots&x_{1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{N-1}&x_{N-2}&\ldots&x_{N-p}\end{bmatrix}\begin{bmatrix}a_{1}\\ a_{2}\\ \vdots\\ a_{p}\end{bmatrix}=-\begin{bmatrix}x_{p}\\ x_{p+1}\\ \vdots\\ x_{N}\end{bmatrix}\hskip 28.45274pt\implies X\cdot a=-x,$ (24)
$\omega_{j}=i\frac{\ln(z_{j})}{h}.$ (25)
Having computed the quasi-normal frequencies, we can also obtain their corresponding amplitudes using the Prony fit. For this, we need to define a $(p\times N)$ matrix $A_{ij}=z_{i-1}^{j}$ and a vector $v=[x_{1},x_{2},\ldots,x_{N}]$. We get the amplitudes by calculating $C=[C_{1},C_{2},\ldots,C_{p}]=(A^{\dagger}A)^{-1}A^{\dagger}v$.
In the method discussed above, we get $p$-number of solutions, but most of them are numerical artefacts and not actual QNMs of the system. To mitigate this issue, we can repeat the above steps with different values of $h$ and $N$ and accept only the results common to both sets of solutions (with appropriate tolerance) as the actual QNMs.
The time-domain and Prony fit method works well for most effective potentials which do not have any discontinuities. But this method often fails to return the $\omega_{QNM}$ with low relative amplitude (we notice this feature for odd overtones in Sec. IV).
III.2 The spectral method
Using the corresponding equation for a perturbation (scalar or axial gravitational), another method known as the spectral method, may also be employed to find the QNMs of the system. In our analysis, we will use the Chebyshev grid and thus we have to map our coordinate to a finite interval, i.e. $[-1,1]$. The tortoise coordinate $l\in(-\infty,+\infty)$ may be mapped to $x\in[-1,1]$ using the transformation $x=\frac{2}{\pi}arctan(l)$ Batic and Dutykh (2024).
$\frac{d^{2}\psi}{dl^{2}}+(\omega^{2}-V(l))\psi=0$ (26)
$\psi(x)=exp(i\omega\frac{2l}{\pi}arctan(l))\phi(x)$
$\phi(x)\approx\sum_{i=0}^{N}\phi(x_{i})C_{i}(x)$ (27)
with $x_{i}$ being the collocation points (Chebyshev roots) given by $x_{i}=\frac{(2i+1)\pi}{2(N+1)}$ (i = 0,1,2,…,N).
Thus, we can express the master equation in the form

### v2

@@ Abstract
Finally, we show how an effect somewhat similar to the so-called ‘environment induced redshift’ of QNMs, introduced recently (Phys. Rev. D 111, 064026 (2025)), may be obtained for spacetimes with non-constant $g_{00}$, via an appropriate tuning of available metric parameters which systematically modify the shapes of the effective potentials arising in the perturbation equations.
@@ I Introduction and summary
The entire scenario was largely theoretical, with very little effort being given to even suggest possible observational consequences. A large part of the efforts focused on how to avoid energy condition violations Visser et al. (2003); Maeda and Nozawa (2008), and modified/alternate theories of gravity Bhawal and Kar (1992); Kanti et al. (2012); Hochberg (1990); Agnese and La Camera (1995); Shaikh and Kar (2016); Kumar et al. (2024); Radhakrishnan et al. (2024); Dehghani and Dayyani (2009); Sadatian and Hosseini (2024); Loewer et al. (2024); Battista et al. (2024) seemed to be a sort of way out. It was noticed that in a modified theory of gravity, the purely geometric violation of the timelike or null convergence condition does not necessarily imply a violation of the weak, null or dominant energy conditions. Hence, energy conditions, postulated as physical requirements on matter, may hold even though the convergence conditions may be violated, as required for a wormhole to exist. Thus, wormholes seemed more viable in various modified theories of gravity.
@@
Given the above fact that wormholes could be around if any modified theory of gravity is acceptable, even in a restricted domain of the theory parameters, it is obvious that one must find observational signatures. Gravitational lensing of wormholes Cramer et al. (1995); Dey and Sen (2008); Shaikh et al. (2019); Godani and Samanta (2021); Nandi et al. (2006); Eiroa et al. (2001); Izmailov et al. (2019); Mitra and Ghosh (2024); Tsukamoto and Harada (2017) was one of the first among the observationally relevant works pursued by researchers. It still continues as an important possible signature.
@@
Over the last decade or so, astronomy and astrophysics has reached a different level of observational relevance through the emergence of a large amount of data and results in gravitational wave astronomy (LIGO-VIRGO-KAGRA collaboration Abbott and others (2023)) and infrared imaging of astrophysical objects (the Event Horizon Telescope (EHT) Akiyama and others (2019, 2022)). The results obtained tally well with General Relativity predictions, leaving little room for modified gravity. However, there does remain a small but definite possibility of accommodating a modified gravity solution or a sort of exoticity within GR (eg. the Bronnikov-Ellis geometry). It is this possibility which has driven some researchers (including us) to construct models, do calculations with the aim of looking for signatures of wormholes in available data. With more observed events and data in the future, one can always hope for further confirmation of GR or, perhaps, some acceptance of a modification of GR, of course, within a specific domain of parameter values.
@@
One of the central observationally important arenas in GW astrophysics, which has been flooded, in the last decade, by a huge amount of theoretical work by various groups, is the study of quasinormal modes Vishveshwara (1970) in perturbation theory (see Berti et al. (2025) for a very recent review). Scalar and gravitational perturbations have been worked out for diverse black hole Berti ; Berti et al. (2009); Cardoso and Lemos (2001); Nollert (1999); Konoplya (2002, 2003); Rincón and Santos (2020); Blázquez-Salcedo et al. (2018b) and wormhole Bueno et al. (2018); Konoplya and Molina (2005); Kim (2008); Alfaro et al. (2024); Jirón et al. (2025) spacetimes in GR as well as in modified gravity theories. Any such study on a spacetime not among the conventional solutions of GR (eg. Schwarzschild or Kerr) is referred to as a study on a ‘mimicker’. More specifically, wormholes constitute a prominent family among such ‘black hole mimickers’ Abdikamalov et al. (2019); Konoplya et al. (2019); Urbano and Veermäe (2019); Narzilloev et al. (2020); Olivares et al. (2020); Shaikh et al. (2022); Shao et al. (2021); Siemonsen (2024); Pani et al. (2010); Cardoso and Pani (2019); Konoplya and Zhidenko (2016); Glampedakis and Pappas (2018); De Simone et al. (2025). Collectively, such mimickers are also referred to as ‘exotic compact objects’. The main idea here is that these spacetimes could be among possible stable end-states in a post-merger scenario Sanchis-Gual et al. (2019); Cardoso et al. (2016); Krishnendu et al. (2017). Whether this is a feasible option at all, will be evident as more observational results emerge.
@@
$\bullet$ Most of the work on QNMs of the family of spacetimes studied here is focused on the fundamental QNMs, which surely have the largest contribution (amplitude-wise) in a time domain profile. Despite this, with possible detection Capano et al. (2023) (though disputed Cotesta et al. (2022)) of overtones and the possible detectability of overtones in LISA Oshita and Tsuna (2023), it is always important to track down the contributions of the so-called ‘overtones’. This, indeed, has been a line of thought of late, with varied motivations. Researchers were keen on finding overtones from the onset of QNM analysis. Chandrasekhar and Detweiler Chandrasekhar and Detweiler (1975) computed the fundamental mode along with a couple of overtones for the Schwarzschild metric. Later, the QNM spectra with overtones for RN Ferrari and Mashhoon (1984), Kerr Detweiler (1980); Leaver (1985), Kerr-Newman Berti and Kokkotas (2005), regular blackhole Konoplya et al. (2022); Konoplya and Zhidenko (2024); Zhang et al. (2024), parametrized black hole Hirano et al. (2024) metrics were computed. The authors of Kim (2008); Batic and Dutykh (2024); Alfaro et al. (2024); Ávalos and Contreras (2022); Jirón et al. (2025); González et al. (2022); Konoplya and Stashko (2025) computed QNM overtones for various wormhole geometries.
@@
$\bullet$ It has been found (theoretically) that black hole quasinormal modes, when obtained for a scenario where the black hole is in an environment (e.g. a halo), exhibit a ‘redshift’ w.r.t. the increase of the environment parameter Barausse et al. (2014); Pezzella et al. (2025); Liu et al. (2025); Chen and Kotlařík (2023).
We show here, for the QNMs and overtones found by us for the wormhole family, that there is indeed a ‘redshift’ when we modify the so-called effective potential by tuning a metric parameter appropriately.
Our article, which is largely a sequel to Dutta Roy et al. (2020), is organised as follows. In the next section, we provide a summary of the wormhole family we study in this article. In Section III, we summarise the numerical results obtained with explicit plots. Section IV discusses the inferences which may be drawn from our studies on the QNMs and overtones. The above-mentioned ‘redshift’ effect is discussed extensively in Section V. Finally, in Section VI, we conclude with remarks on possible future work. Appendix A quickly outlines the methods adopted in finding the QNMs–more specifically, which method is better suited for finding overtones is emphasised and discussed in more detail here. The appendix B contains the detailed numerical values of the QNMs, thereby supporting the plots shown in Section III.
@@ II A brief review on the family of wormhole geometries
One may also rewrite the line element using the proper radial distance, i.e $r^{2}=b_{0}^{2}+l^{2}$ (where $\infty\leq l\leq\infty$). In this representation, we have
@@
We may visualise the above line element as representing two asymptotically flat regions ($l\rightarrow\pm\infty$) connected by a throat/bridge of radius $b_{0}$. Note that the radial coordinate $r$ runs from $b_{0}$ to $\infty$ whereas $l$ runs from -$\infty$ to $\infty$ with $l=0$ corresponding to $r=b_{0}$.
@@ II.1 The ultra-static wormhole family
Here, $l$ $(\in[-\infty,\infty])$ is the tortoise coordinate
@@ II.2 Bronnikov–Ellis wormhole with asymptotic mass
Till now, we have discussed the ultrastatic, symmetric and massless Bronnikov–Ellis geometry and its generalisation. However, the authors of Bronnikov (1973), Ellis (1973) (see also Blázquez-Salcedo et al. (2018a)) have proposed a metric for a massive, static, Bronnikov–Ellis wormhole for which the line element is of the following form:
@@
$f(l)=\frac{C}{b_{0}}\left(\arctan\left(\frac{l}{b_{0}}\right)-\frac{\pi}{2}\right)$
@@
For $M=0$, we recover the symmetric generalised Bronnikov–Ellis wormhole, while any non-zero value of $M$ will introduce asymmetry in the wormhole shape. The QNMs for the $n=2$, $M\neq 0$ geometry will be discussed in Section V.1.
@@
However, in the case of $M\neq 0$ and $n>2$, one can find that all three conditions are satisfied in certain regions if the $|M|$ exceeds a specific value (Fig. 1). For different geometries (different values of $n$), the critical value of $M$ and the acceptable regions (corresponding to a fixed value of $M$) are different.
@@ II.3 Bronnikov–Ellis geometry with a chosen redshift function:
It is also possible to choose a suitable redshift function ($g_{00}$ never zero) in an ad-hoc way while retaining the Bronnikov–Ellis $n=2$ character unaltered in the spacelike slice. The ensuing consequences for required matter and energy conditions can thereafter be looked at. As an example, let us consider the line element studied in Alfaro et al. (2024) and given by,
@@
Note that the energy density is not affected by $M$, though the pressures change. The Weak Energy Condition is violated for all $r>b_{0}$ since the energy density is negative. The Null Energy Condition is also violated with complementary behaviour for the $\rho+\tau$ and $\rho+p$ inequalities w.r.t. increase in $M$ value (Fig. 2). As $M$ increases the region of violation (near $r=b_{0}$) of the $\rho+\tau$ inequality shrinks whereas the region of violation (away from $r=b_{0}$) of the $\rho+p$ inequality becomes larger. We have plotted $\rho+\tau$ and $\rho+p$ as a function of the radial coordinate $r$, for two values of $M$ (Fig. 2), in order to show this feature in the NEC violations.
Putting $M=0$, we get back the original Bronnikov–Ellis wormhole. We will discuss the effective potential and the $\omega_{QNM}$ for this spacetime in a later section (V.2).
@@ II.4 Effective potentials for scalar and axial perturbation
The original Bronnikov–Ellis wormhole ($n=2$) is a solution to the Einstein equations coupled to a phantom scalar field. The $n>2$ geometries also have contributions from the phantom field. The perturbations in this phantom scalar field ($\Phi$) can give rise to scalar quasinormal modes. For this, we may write the scalar field as $\Phi+\phi$. Here $\phi$ is the perturbation on the phantom scalar field and it satisfies
@@
$\frac{d^{2}\psi}{dl_{*}^{2}}+(\omega^{2}-V_{sc}(l))\psi=0\hskip 56.9055pt\text{where}\hskip 28.45274pt\frac{dl_{*}}{dl}=e^{-f}$ (16)
@@
III Results: quasinormal spectra (fundamental and overtones for the ultra-static case)
The problem of computing quasi-normal modes for the Bronnikov–Ellis wormhole has been tackled in many previous works. The authors of Kim (2008); Blázquez-Salcedo et al. (2018a) and Dutta Roy et al. (2020); Dutta Roy (2022) have presented the fundamental modes for the original Bronnikov–Ellis ($n=2$ case) and the generalised Bronnikov–Ellis wormholes, respectively.